OpenStudy (anonymous):

What is the distance between point G and point H? Round your answer to the nearest tenth. A. 3.2 units B. 4.3 units C. 5.8 units D. 7.1 units

OpenStudy (anonymous):

use distance formula

OpenStudy (anonymous):

I don't know what that is!!! That is why I am on here!!!

OpenStudy (anonymous):

i showed on the last problem http://www.purplemath.com/modules/distform.htm

OpenStudy (calculusxy):

do you get it or still confused?

OpenStudy (anonymous):

Still confused, sorry, all these number just get jumble up in my head.

OpenStudy (calculusxy):

no problem. when you connect the vertices for the points, you will see that it a triangle.

OpenStudy (anonymous):

OK.....

OpenStudy (calculusxy):

do you know the Pythagorean theorem?

OpenStudy (anonymous):

Roughly, I am really bad at that.

OpenStudy (calculusxy):

it is:

OpenStudy (calculusxy):

a^2+b^2=c^2 c^2 is the hypotenuse (longest side of the triangle)

OpenStudy (anonymous):

ok so A x A + B x B = C x C ?

OpenStudy (calculusxy):

yes. but on c^2, you want to find the square root to find the actual value of c (since c is squared)

OpenStudy (anonymous):

How do you sqaure C, wait what is C?

OpenStudy (calculusxy):

hypotenuse

OpenStudy (calculusxy):

my suggestion is to analyze the picture that I gave you. this is my work in visual.

OpenStudy (anonymous):

So C is a right Triangle?

OpenStudy (calculusxy):

the whole triangle is the right triangle, C is the longest side of the triangle (hypotenuse)

OpenStudy (anonymous):

So the answer is C?

OpenStudy (calculusxy):

yes. do you have any other confusions? you can tell me because I can help you for further reference.

OpenStudy (anonymous):

noPe :D THANKS!!

OpenStudy (calculusxy):

welcome